OpenStudy (anonymous):

Which equation is a quadratic equation? 3x − y = 8x + 2 y = 7x(5 − 2x) + 9 7x2 + y = (3 + x)(7x − 1) y − 3 = (x2 − 3)(8x + 3)

OpenStudy (anonymous):

A quadratic equation has an x^2 term in it. The x^2 term is the largest exponent. 3x - y = 8x + 2 -y = -3x + 8x + 2 -y = 5x + 2 y = -5x - 2 This is not a quadratic equation.

OpenStudy (anonymous):

y = 7x(5 - 2x) + 9 y = 35x - 14x^2 + 9 y = -14x^2 + 35x + 9 This is a quadratic equation.

OpenStudy (anonymous):

7x^2 + y = (3 + x)(7x - 1) 7x^2 + y = 21x - 3 + 7x^2 - x 7x^2 + y = 20x - 3 + 7x^2 y = 20x - 3 + 7x^2 - 7x^2 y = 20x - 3 Even though this equation looks like it has an x^2 term, after solving for y, you can see that it is not a quadratic equation.

OpenStudy (anonymous):

No you try the last one.

OpenStudy (anonymous):

*Now

OpenStudy (anonymous):

@Aaron0620, Do you know how to solve for y in the last equation?

OpenStudy (anonymous):

urgg fine i will lol, and yes ill try

OpenStudy (anonymous):

Start by multiplying the two binomials together.

OpenStudy (anonymous):

7x^2 + y = (3 + x)(7x − 1) -7x^2 -7x^2 -------------------------- y= -7x^2+20x -3 is that right?

OpenStudy (anonymous):

7x^2 + y = 21x - 3 + 7x^2 - x 7x^2 + y = 20x + 7x^2 - 3 y = 20x + 7x^2 - 3 - 7x^2 y = 20x - 3 You are close, but the 7x^2's cancel each other out. So even though it looks like there is an x^2 term in the beginning, after you solve for y, you can see that it is not a quadratic equation. The only quadratic equation is #2. You are welcome. :)

OpenStudy (anonymous):

alright thank you!

OpenStudy (anonymous):

No problem. c: